OpenStudy (anonymous):

what is-6squared+-6x5+5squared

OpenStudy (anonymous):

-41

OpenStudy (anonymous):

is tht right?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

ok thnx bro

OpenStudy (anonymous):

You may click good answer if you approve

OpenStudy (anonymous):

It helps

OpenStudy (anonymous):

ok i did

OpenStudy (anonymous):

ur dumb thts wrong my teacher just did it

OpenStudy (anonymous):

look for yourself

OpenStudy (anonymous):

srry my teacher did it wrong and ur right bro sorry

OpenStudy (anonymous):

:) No worries.

OpenStudy (anonymous):

your a beast

OpenStudy (anonymous):

ur a beast

OpenStudy (anonymous):

I am not up to date on internet lingo...lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-6squared+-6x5+5squared I assume: \[-6^{2} - 30 + 25 = - 36 - 5 = - 41\]